# NeRF-CA: Dynamic Reconstruction of X-ray Coronary Angiography with Extremely Sparse-views

### v1

@@ Abstract
Dynamic three-dimensional (4D) reconstruction from two-dimensional X-ray coronary angiography (CA) remains a significant clinical problem. Challenges include sparse-view settings, intra-scan motion, and complex vessel morphology such as structure sparsity and background occlusion. Existing CA reconstruction methods often require extensive user interaction or large training datasets. On the other hand, Neural Radiance Field (NeRF), a promising deep learning technique, has successfully reconstructed high-fidelity static scenes for natural and medical scenes. Recent work, however, identified that sparse-views, background occlusion, and dynamics still pose a challenge when applying NeRF in the X-ray angiography context. Meanwhile, many successful works for natural scenes propose regularization for sparse-view reconstruction or scene decomposition to handle dynamics. However, these techniques do not directly translate to the CA context, where both challenges and background occlusion are significant. This paper introduces NeRF-CA, the first step toward a 4D CA reconstruction method that achieves reconstructions from sparse coronary angiograms with cardiac motion. We leverage the motion of the coronary artery to decouple the scene into a dynamic coronary artery component and static background. We combine this scene decomposition with tailored regularization techniques. These techniques enforce the separation of the coronary artery from the background by enforcing dynamic structure sparsity and scene smoothness. By uniquely combining these approaches, we achieve 4D reconstructions from as few as four angiogram sequences. This setting aligns with clinical workflows while outperforming state-of-the-art X-ray sparse-view NeRF reconstruction techniques. We validate our approach quantitatively and qualitatively using 4D phantom datasets and ablation studies.
@@
X-ray coronary angiography, 4D reconstruction, Neural Radiance Field, sparse-view, 3D reconstruction
@@ I Introduction
X-ray is fundamentally limited by providing a two-dimensional (2D) representation of a three-dimensional (3D) structure, resulting in a lack of depth perception. This problem is especially apparent in X-ray coronary angiography (CA) interventions, where complex dynamic 3D (4D) vessel structures are represented by dynamic sequences of 2D X-ray projections or coronary angiograms. On top of that, in standard CA interventions, an extremely limited set of these angiogram views are acquired per patient, which we refer to as sparse-view [1]. These interventions could benefit from 4D reconstructions, for example, to enhance the 4D perception of the vessel structure for catheter navigation applications [2]. Nonetheless, accurate 4D reconstruction from CA data poses a challenge due to inherent characteristics from the X-ray system, complex vessel characteristics, and intra-scan motion [3]. The challenges of the X-ray system include extremely sparse-view X-ray projections and system geometry inaccuracies. The vessel characteristics include structure sparsity, vessel overlap, and background occlusion due to overlapping background structures or contrast inhomogeneity. Moreover, coronary angiography data exhibits intra-scan motion due to cardiac and respiratory motion. Although clinically relevant, the 4D reconstruction of CA data remains challenging.
@@
Several CA reconstruction methods have been proposed, including traditional and machine learning methods [3, 4]. Both methods can obtain satisfactory reconstructions from extremely sparse-view sequences. However, they rely on time-consuming user interactions, error-prone background removal techniques, or large amounts of training data. These dependencies hinder the adaptation of these methods in clinical practice. While some methods ignore the dynamics, others assume cyclic cardiac motion through a simultaneously acquired electrocardiogram (ECG) signal but still lead to motion artifacts [3].
@@
Neural Radiance Field (NeRF) is a promising deep learning technique for 3D reconstruction that can reconstruct high-fidelity 3D natural scenes from 2D projections [5]. Unlike machine learning methods that learn a prior, NeRF learns an individual scene given a set of 2D images at hand. As such, it does not rely on segmentations, human interaction, or large amounts of training data, showing the potential to overcome the limitations of prior CA reconstruction techniques. Recent work explored the potential of NeRF in X-ray angiography, showing that NeRF limitations still lie in reconstruction in a sparse-view setting [6]. Meanwhile, many extended works of NeRF have been proposed to reconstruct in the sparse-view setting [7]. Regularization-based techniques for natural scenes have been shown to reconstruct natural scenes from only $3$ views [8]. However, techniques proposed for medical scenes do not reach the extremely sparse-view requirements for CA scenes. Natural-scene-based techniques do not directly apply to medical scenes, as they rely on scene-specific properties, such as depth or smoothness, that do not occur in medical scenes. Furthermore, methods for reconstructing video-acquired natural dynamic scenes have also been proposed [9]. Liu et al. [10] recently presented a technique consisting of static and dynamic scene decomposition for modeling vessel dynamics in digital subtraction angiography (DSA) separating foreground and background. Although similar to CA, DSA does not suffer from background occlusion, and the technique proposed by Liu et al. is not adequate for sparse views since requires many training projections.
@@
We propose NeRF-CA, the first step toward a 4D X-ray Coronary Angiography (CA) reconstruction method addressing sparse-views and the cardiac motion characteristics of CA acquisitions. We present 4D vessel reconstructions from an extremely sparse amount of coronary angiogram sequences exhibiting cardiac motion. We leverage the extended work of NeRF for static and dynamic natural scene decomposition [11] and techniques for reconstruction in a sparse-view setting [8, 12]. Specifically, we utilize the rapid motion of the coronary artery to decouple the scene into a dynamic coronary artery component and a static background component. To perform the dynamic reconstruction, we uniquely combine scene decomposition with a weighted pixel sampling technique, static and dynamic factorization regularization, and entropy regularization, enforcing an accurate foreground and background separation. For sparse-view reconstruction, we combine these techniques with coarse-to-fine positional encoding and occlusion regularization. Together, we demonstrate vessel reconstruction capabilities from as few as four coronary angiogram sequences, matching the clinical workflows. We validate the relevance of these constraints through ablation studies. We demonstrate our method on two 4D phantom datasets [13, 14], simulating synthetic coronary angiograms. These phantom datasets allow us to perform a quantitative validation, as we can acquire ground-truth sequences from any viewpoint. We also compare our method to X-ray sparse-view NeRF reconstruction methods, significantly outperforming them in the extremely sparse-view X-ray coronary angiography setting.
@@
In summary, the main contribution of our work is a novel method for 4D reconstruction of extremely sparse-view CA addressing Unlike prior methods, our method does not rely on manual user interactions or large datasets. We decouple the coronary angiography scene in a dynamic coronary artery and static background by uniquely combining scene decomposition with regularization constraints fitting the sparse nature of the blood vessel structure. Moreover, we perform experiments to compare our method to other methods while also validating the relevance of our choices through ablation studies.
@@ II-A X-ray Coronary Angiography
X-ray coronary angiography (CA) is the most common imaging modality to diagnose and treat coronary artery disease. Even though there are fundamental limitations due to its 2D representation of a 3D anatomy, it still outweighs 3D imaging modalities. The benefits are its usage for diagnosis and treatment and spatial and temporal resolution [15]. This work focuses on CA imaging acquired by a single-plane C-arm angiography system, which is still the gold standard for CA interventions [16].
@@
A coronary angiogram sequence consists of multiple frames displaying a contrast-enhanced coronary artery experiencing cardiac and respiratory motion. Figure 1 provides an example of one coronary angiogram sequence generated from two synthetic datasets across multiple cardiac phases $i$ within one cardiac cycle. As can be seen, the imaging contains sparse blood vessel structures that may overlap. Moreover, background occlusion occurs, either through overlapping background structures or contrast inhomogeneity. In a clinical setting, an extremely sparse amount of angiogram sequences (i.e., 2-4 per patient) is captured to minimize X-ray exposure. C-arm systems can be rotated using 3D Euler angles $\theta$ and $\phi$ to obtain varying viewing points [3]. The operator chooses these viewing points, typically based on standard optimal viewing points described in literature [17]. The coronary arteries experience cardiac and respiratory motion while acquiring the angiogram sequences [18]. The cardiac motion of the coronary arteries aligns with the beating motion of the heart, which is typically assumed to be synchronous with the simultaneously recorded electrocardiogram (ECG). Therefore, individual frames of the coronary angiogram can be assigned an estimated cardiac phase within each cardiac cycle, as shown in Figure 1. Cardiac motion dominates the dynamics [18]. However, respiratory motion also occurs, which is correlated with the breathing rate.
4D reconstructions in the context of CA can aid both the diagnosis and treatment of coronary artery disease. The 2D CA sequences may lead to a sub-optimal assessment of stenosis severity and, therefore, sub-optimal stent size selection, which a 4D reconstruction could overcome. Coronary roadmaps are commonly generated from a reconstruction to aid catheter navigation through the blood vessels[2]. The reconstruction quality for this roadmap application is typically evaluated based on the vessel topology rather than the vessel diameter accuracy [6].
@@ II-B Neural Radiance Field
A Neural Radiance Field (NeRF) represents a single scene through a multilayer perceptron (MLP), a fully connected neural network [5]. Given a 3D coordinate $\boldsymbol{x}=(x,y,z)$ and viewing direction $\boldsymbol{d}=(\theta,\phi)$, the network outputs a density $\sigma$ and color $\boldsymbol{c}=(r,g,b)$ to represent natural scenes. NeRF therefore learns a continuous function $F_{\Theta}(\boldsymbol{x},\boldsymbol{d})=(\sigma,\boldsymbol{c})$, where $\Theta$ represents the learned weights of the MLP.
As MLPs have difficulty learning high-frequencies, NeRF employs positional encoding $\gamma$ on the input parameters to allow for high-frequency details in the reconstructions [5]. The earliest works used sinusoidal functions with different frequencies to map the input to a higher dimensional space [19].
NeRF renders pixels through the learned density $\sigma$ and color $\boldsymbol{c}$ by discretizing the volume rendering equation [20]. Every pixel is associated with a ray $\boldsymbol{r}$, which is defined as $\boldsymbol{r}(t)=\boldsymbol{o}+t\boldsymbol{d}$, with camera origin $\boldsymbol{o}$, ray direction $\boldsymbol{d}$ and $t$ the distance from the origin. The obtained predicted pixel color $\hat{C}(\boldsymbol{r})$ is compared to the ground truth pixel color $C(\boldsymbol{r})$ of a given image to optimize the model utilizing mean squared error (MSE). We refer to the work of Mildenhall et al. [5] for further details on the method.
@@ III-A X-ray coronary angiography reconstruction techniques
A wide range of X-ray coronary angiography (CA) reconstruction techniques have been proposed, which can be categorized into model-based, tomographic-based, and machine learning-based [3, 4]. We will discuss these techniques, their background removal strategies, and their strategies to handle cardiac motion.
Model-based techniques extract the vessels from the CA imaging to reconstruct them as binary trees. These methods achieve accurate reconstruction with sparse clinical views but rely on time-consuming manual input [15]. Recently, machine learning approaches have been proposed for automatic segmentation [21, 22] or direct 3D reconstruction [4, 23]. However, these methods require extensive segmented training data, often unavailable in coronary angiography. Tomographic techniques reconstruct attenuation volumes from raw X-ray imaging without needing manual segmentation. However, they require many views, limiting their use in single-plane CA. Methods based on NeRF could potentially overcome these limitations, which will be discussed in Section III-B. Our work will build on these methods in the context of CA.
CA imaging captures both coronary arteries and background structures, which can cause truncation errors due to incomplete visibility from every viewpoint. As a result, background removal through suppression or subtraction is a common technique applied in tomographic reconstruction. As a pre-processing step, good suppression and subtractions can be achieved, but they are hindered by the dynamics of the CA imaging, leading to background motion artifacts [24]. Other works perform background subtraction during reconstruction, which have shown to be effective in separating te vessel from the background [25]. Inspired by these methods, our work also separates the blood vessels from the background to avoid truncation errors by decoupling the dynamic coronary artery from the static background.
Traditional 3D reconstruction methods use various strategies to mitigate motion artifacts, primarily focusing on cardiac motion and treating respiratory motion as residual [3]. They use ECG signals for frame extraction through gating or motion compensation. Gating selects frames around the same cardiac phase, while motion compensation aligns all frames to the same phase. Our work focuses on cardiac motion, assuming residual respiratory motion. Incorporating respiratory motion should be explored in future work.
@@ III-B Neural Radiance Fields
NeRFs have been widely adopted for reconstructing natural scenes due to their ability to reconstruct high-fidelity scenes. However, challenges persist in reconstructing natural scenes from sparse-views, especially when dynamic [26]. On top of that, medical imaging scenes present unique challenges such as sparse structures and background occlusion [27, 28].
@@
Many works have addressed the sparse-view problem, which can be categorized into prior-based and regularization-based methods [7]. Whereas prior-based methods train a general model on large datasets, regularization-based methods introduce additional scene-specific regularization terms. Given the lack of data in single-plane CA, as discussed in Section III-A, prior-based methods are unsuitable in our context. Regularization-based methods instead rely on scene-specific properties [29, 30, 12, 8]. However, these methods are yet to be adapted to medical scenes, which pose unique properties such as background occlusion. For medical scenes, reconstruction in a sparse-view setting also poses a significant challenge [31, 32]. Prior-based [33, 34, 35] or X-ray regularization-based [36, 37, 32] strategies have been proposed to address this in static scenes, reaching a limit in performance with $15$ projections [36]. However, this number of projections is still inadequate for our clinical context and does not account for sparse structures in dynamic scenes. In our work, we leverage the regularization-based methods proposed for natural scenes by adapting them to the blood vessel properties.
@@
On the other hand, many strategies have been proposed for the 4D reconstruction of dynamic scenes [9]. For medical scenes, some works have been proposed to model the dynamics, but they require a significantly large amount of projections [32, 38, 39]. NeRF methods have also been proposed for blood vessel reconstruction, addressing sparse-views and dynamics. For vessels that are not prone to extensive motion, digital subtraction angiography (DSA) can be used to remove all background structures from the 2D projection images. TiaVox [40] and Liu et al. [10] successfully reconstruct these sparse structure scenes from only $30$ DSA projections. Liu et al. [10] model blood flow dynamics by decoupling the scene into a static background and dynamic blood vessel foreground. Static and dynamic scene decomposition is one of the approaches to model dynamics with NeRFs, where scenes can be decoupled in a static and dynamic component [26]. However, these DSA scenes do not experience background occlusion, recently highlighted as a significant challenge in sparse-view X-ray angiography reconstruction with NeRFs [6]. Moreover, CA scenes exhibit even more significant dynamics in foreground and background than DSA scenes. In this work, we decouple the static and dynamic scenes and extend this strategy to address the properties of CA scenes. We build on a natural scene decomposition technique D2NeRF [11], which we adjust to CA scenes. On top of that, we impose regularization on the individual decoupled scenes, including structure sparsity [12], scene smoothness, and edge occlusion minimization [8]. These approaches are the bases of NeRF-CA, which allows for 4D reconstruction from an extremely sparse number of angiogram sequences.

### v2

@@ Abstract
Dynamic three-dimensional (4D) reconstruction from two-dimensional X-ray coronary angiography (CA) remains a significant clinical problem. Existing CA reconstruction methods often require extensive user interaction or large training datasets. Recently, Neural Radiance Field (NeRF) has successfully reconstructed high-fidelity scenes in natural and medical contexts without these requirements. However, challenges such as sparse-views, intra-scan motion, and complex vessel morphology hinder its direct application to CA data. We introduce NeRF-CA, a first step toward a fully automatic 4D CA reconstruction that achieves reconstructions from sparse coronary angiograms. To the best of our knowledge, we are the first to address the challenges of sparse-views and cardiac motion by decoupling the scene into the moving coronary artery and the static background, effectively translating the problem of motion into a strength. NeRF-CA serves as a first stepping stone for solving the 4D CA reconstruction problem, achieving adequate 4D reconstructions from as few as four angiograms, as required by clinical practice, while significantly outperforming state-of-the-art sparse-view X-ray NeRF. We validate our approach quantitatively and qualitatively using representative 4D phantom datasets and ablation studies. To accelerate research in this domain, we made our codebase public: https://github.com/kirstenmaas/NeRF-CA.
@@
4D reconstruction, X-ray coronary angiography, Neural Radiance Field, sparse-view, static and dynamic decomposition
@@ I Introduction
X-rays are fundamentally limited by providing a two-dimensional (2D) representation of a three-dimensional (3D) structure, resulting in a lack of depth perception. This problem is especially apparent in X-ray coronary angiography (CA) interventions, where dynamic sequences of 2D X-ray projections or coronary angiograms represent complex dynamic 3D (4D) vessel structures. On top of that, in standard CA interventions, an extremely limited set of these angiogram views are acquired per patient, which we refer to as sparse-view [1]. These interventions could benefit from 4D reconstructions to enhance the 4D perception of the vessel structure for catheter navigation applications [2]. Nonetheless, accurate 4D reconstruction from CA data poses a challenge due to inherent characteristics from the X-ray system, complex vessel characteristics, and intra-scan motion [3]. The challenges of the X-ray system include extremely sparse-view X-ray projections and system geometry inaccuracies. The vessel characteristics include structure sparsity, vessel overlap, and poor vessel visibility due to background obstruction, related to overlapping background structures or blood vessel contrast inhomogeneity. Moreover, coronary angiography data exhibit intra-scan motion due to cardiac and respiratory motion. Overall, 4D reconstruction of CA data remains a challenge.
@@
Several CA reconstruction methods have been proposed, including traditional and machine learning methods [3, 4, 5]. Both methods can obtain satisfactory reconstructions from extremely sparse-view sequences. However, they rely on time-consuming user segmentations or large amounts of training data, consisting of many views per patient, which are not available due to the clinical sparse-view setting of CA. These dependencies hinder the adaptation of these methods in clinical practice. While some methods ignore the dynamics, others assume cyclic cardiac motion through an acquired electrocardiogram (ECG) signal but still lead to motion artifacts [3].
@@
Neural Radiance Field (NeRF) is a promising deep learning technique for 3D reconstruction that can reconstruct high-fidelity 3D natural scenes from 2D projections [6]. Unlike machine learning methods that learn a prior, NeRF learns an individual scene given a set of 2D images at hand. As such, it does not rely on segmentations or large amounts of training data, showing the potential to overcome the limitations of prior CA reconstruction techniques. Recent work explored the potential of NeRF in X-ray angiography, showing that limitations still lie in reconstruction in a sparse-view setting [7].
Meanwhile, many extended NeRF works have been proposed to reconstruct in the sparse-view setting [8]. Regularization-based techniques for natural scenes have been shown to reconstruct from only $3$ views [9]. Techniques designed for natural scenes rely on properties such as depth and sparsity, which are absent in the CA setting, making them unsuitable. Concurrently, NeRF-based techniques proposed for medical scenes do not reach the extremely sparse view requirements for CA scenes, requiring a minimum of 15 projections for adequate performance [10]. NeRF methods, such as 3DGR-CAR [11] and NeCA [12], address sparse-views for CA scenes. However, they rely on accurate segmentation or extensive training data and ignore the inherent motion of CA data. On the other hand, Liu et al. [13] recently presented static and dynamic scene decomposition for modeling vessel dynamics in digital subtraction angiography (DSA) separating foreground and background. Although similar to CA, DSA does not suffer from poor vessel visibility, and the technique proposed by Liu et al. [13] is not adequate for sparse-view settings since it requires many training projections.
@@
We propose NeRF-CA, the first step towards a 4D X-ray Coronary Angiography (CA) reconstruction method addressing the challenges of sparse-views and the cardiac motion characteristics of CA acquisitions. We present 4D vessel reconstructions from an extremely sparse amount of synthetic coronary angiogram sequences exhibiting cardiac motion. We build on the extended work of NeRF for static and dynamic natural scene decomposition [14] and techniques for reconstruction in a sparse-view setting [9, 15]. Specifically, we utilize the rapid motion of the coronary artery to decouple the scene into a dynamic coronary artery component and a static background component. The novelty of our work lies in developing a combination of self-supervised decomposition with regularization techniques to enable sparse-view reconstructions for CA scenes. Specifically, we propose regularization techniques for the dynamic coronary artery component, enforcing properties such as scene sparsity and minimal occlusion. Together, we demonstrate vessel reconstruction capabilities from as few as four coronary angiogram views, matched to clinical workflows. We validate the effect of these constraints through ablation studies. Similar to other related works [4, 11], we demonstrate our method on two 4D phantom datasets [16, 17], simulating synthetic coronary angiograms. These synthetic coronary angiograms allow us to address the challenges of sparse-views and cardiac motion in CA data, having a ground-truth as realistic as possible. Meanwhile, we can perform the necessary quantitative validation, as we can acquire ground-truth sequences from any viewpoint. We compare our method with X-ray sparse-view NeRF reconstruction methods, significantly outperforming them in the extremely sparse-view X-ray coronary angiography setting. This provides the initial validation of the method in an ideal setting, showing the potential for future developments and applications to real data.
@@
In summary, the main contribution is a methodology to address cardiac motion and sparse-view characteristics for 4D reconstruction of CA data. This serves as a first step towards a solution for clinical data. Unlike previous methods, our method does not rely on manual user interactions or large datasets. We decouple the coronary angiography scene in a dynamic coronary artery and static background by uniquely combining scene decomposition with regularization constraints fitting the sparse nature of the blood vessel structure. Moreover, we perform experiments to compare our method to existing state-of-the-art NeRF-based methods while validating the relevance of our choices through ablation studies.
@@ II-A X-ray Coronary Angiography
X-ray coronary angiography (CA) is the most common imaging modality to diagnose and treat coronary artery disease. Although it has fundamental limitations due to its 2D representation of a 3D anatomy, it still outperforms 3D imaging modalities. The benefits are its spatial and temporal resolution [18]. This work focuses on CA imaging acquired by a single-plane C-arm angiography system, which is still the gold standard for CA interventions [19].
@@
A coronary angiogram sequence consists of multiple frames showing a contrast-enhanced coronary artery experiencing cardiac and respiratory motion. In single-plane CA, these sequences are obtained from a single acquisition angle or view. The terms ”sequence” and ”view” will be used interchangeably throughout this paper, depending on whether the focus is on the temporal or directional aspect, respectively. In Figure 1, an example of one coronary angiogram sequence is shown for two datasets, where each frame represents a single cardiac phase $i$ within one cardiac cycle. As can be seen, the imaging contains sparse blood vessel structures that may overlap, as highlighted in Figure 1 (a). Moreover, the imaging suffers from poor vessel visibility through overlapping background structures or inhomogeneous contrast, which is highlighted in Figure 1 (b). In a clinical setting, an extremely sparse amount of angiogram views (i.e., 2-4 per patient) is captured to minimize X-ray exposure. C-arm systems can be rotated using 3D Euler angles $\theta$ and $\phi$ to obtain varying viewing points [3]. The coronary arteries experience cardiac and respiratory motion while acquiring the angiogram sequences. Cardiac motion dominates the dynamics of the coronary arteries [20]. This motion aligns with the beating of the heart, which is assumed to be synchronous with the simultaneously recorded electrocardiogram (ECG). Therefore, individual frames of the coronary angiogram can be assigned an estimated cardiac phase within each cardiac cycle, as shown in Figure 1. On the other hand, respiratory motion is correlated with the breathing rate, which is typically not measured clinically. As a result, this motion is more complex to estimate [20] and is assumed to be residual for 3D reconstruction tasks [3].
4D reconstructions in the context of CA can aid both the diagnosis and treatment of coronary artery disease. An example of a clinical application of 4D reconstructions is called coronary roadmap, in which the reconstruction is overlayed over the X-ray to aid catheter navigation through the blood vessels [2]. This overlay is created by re-projecting the reconstruction based on the C-arm angle of interest, similar to the novel view application of NeRFs.
@@ II-B Neural Radiance Field
A Neural Radiance Field (NeRF) represents a single scene through a multilayer perceptron (MLP), a fully connected neural network [6]. Given a 3D coordinate $\boldsymbol{x}=(x,y,z)$ and a viewing direction $\boldsymbol{d}=(\theta,\phi)$, the network outputs a density $\sigma$ and color $\boldsymbol{c}=(r,g,b)$ to represent natural scenes. NeRF therefore learns a continuous function $F_{\Theta}(\boldsymbol{x},\boldsymbol{d})=(\sigma,\boldsymbol{c})$, where $\Theta$ represents the learned weights of the MLP. A NeRF needs to be re-optimized for each scene.
As MLPs have difficulty learning high frequencies, NeRF employs positional encoding $\gamma$ on the input parameters to allow for high-frequency details in the reconstructions [6]. Early works used sinusoidal functions with different frequencies to map the input to a higher dimensional space [21].
NeRF renders pixels through the learned density $\sigma$ and color $\boldsymbol{c}$ by discretizing the volume rendering equation [22]. Every pixel is associated with a ray $\boldsymbol{r}$, which is defined as $\boldsymbol{r}(t)=\boldsymbol{o}+t\boldsymbol{d}$, with camera origin $\boldsymbol{o}$, ray direction $\boldsymbol{d}$ and $t$ the position along the ray. The obtained predicted pixel color $\hat{C}(\boldsymbol{r})$ is compared to the ground truth pixel color $C(\boldsymbol{r})$ of a given image to optimize the model utilizing mean squared error (MSE) through the photometric loss. We refer to the work of Mildenhall et al. [6] for further details on the method.
@@ III-A X-ray coronary angiography reconstruction techniques
A wide range of X-ray coronary angiography (CA) reconstruction techniques have been proposed, which can be categorized into model-based, tomographic-based, and machine learning-based [3, 4]. We will discuss these techniques and their strategies to handle cardiac motion and sparse-views.
Model-based techniques extract the vessels from the CA imaging to reconstruct them as binary trees. These methods achieve accurate reconstruction with sparse clinical views but rely on time-consuming manual input [18]. Recently, machine learning approaches have been proposed for automatic segmentation [23, 24] or direct 3D reconstruction [4, 25, 5]. However, these methods require large amounts of accurately segmented training data, including multiple views per patient, which is unfeasible in clinical settings due to the need to minimize X-ray exposure and the time-consuming nature of segmentation. Although automatically obtained segmentations offer a potential solution, 3D reconstruction methods that depend on them are highly sensitive to segmentation quality. As stated by several works [23, 24, 5], inaccuracies, such as missing vessels or inconsistencies across 2D views, can propagate through the reconstruction process, resulting in noisy or incomplete 3D reconstructions. Tomographic techniques reconstruct attenuation volumes from raw X-ray imaging without needing manual segmentation. However, they require many views, limiting their use in single-plane CA. Methods based on NeRF have the potential to overcome these limitations, and will be discussed in Section III-B. Our work will build on these methods in the context of CA.
Traditional 3D reconstruction methods use various strategies to mitigate motion artifacts, primarily focusing on cardiac motion and treating respiratory motion as residual [3]. For the cardiac motion, they use ECG signals for frame extraction through gating or motion compensation. Gating selects frames around the same cardiac phase, while motion compensation aligns all frames to the same phase.
Our work focuses on solving the challenge of cardiac motion within the sparse-view setting, not considering respiratory motion and other real-world data aspects, like noise or patient motion. NeRF-CA serves as a first step towards a solution for the full clinical problem, given that a method has yet to be proposed that addresses cardiac motion and sparse-views without the reliance on manual segmentations or large datasets, where both factors have so far hindered the development of a clinically applicable solution to the reconstruction problem. Incorporating respiratory motion and other real-world data aspects is out of scope for this paper, such that the cardiac motion and sparse-view challenges can be isolated and evaluated independently. In case of potential results, future work should address the remaining data aspects.
@@ III-B Neural Radiance Fields
NeRFs have been widely adopted for reconstructing natural scenes due to their ability to reconstruct high-fidelity details. However, challenges persist in reconstructing natural scenes from sparse-views, especially when dynamic [26]. On top of that, CA scenes present unique challenges such as sparse structures and poor vessel visibility [27, 28].
@@
Many works have addressed the sparse-view problem, which can be categorized into prior-based and regularization-based methods [8]. Whereas prior-based methods train a general model on large datasets, regularization-based methods introduce additional scene-specific regularization terms. Given the lack of data in single-plane CA, as discussed in Section III-A, prior-based methods are unsuitable in our context. Regularization-based methods instead rely on scene-specific properties [29, 30, 15, 9]. However, these methods are yet to be adapted to medical scenes, which pose unique properties such as poor vessel visibility and therefore complicating foreground and background distinction. For medical scenes, reconstruction in a sparse-view setting also poses a significant challenge [31, 32]. Prior-based [33, 34, 35] or X-ray regularization-based [10, 36, 32] strategies have been proposed to address this in static scenes, reaching a limit in performance with $15$ projections [10]. However, this number of projections is still inadequate for our clinical context and does not account for sparse structures in dynamic scenes. In our work, we are inspired by the regularization-based methods utilized for natural scenes. We propose regularization-based methods that fit blood vessel properties, effectively translating regularization-based techniques to the medical imaging setting.
@@
NeRF methods have also been proposed for blood vessel reconstruction, addressing sparse-views and dynamics independently. For CA scenes, several NeRF-based techniques have been proposed, including 3DGR-CAR [11] and NeCA [12]. These methods address the sparse-view problem, enabling reconstructions from as few as two projections. However, they depend on accurate segmentation or extensive training data and overlook the inherent dynamics of CA data. For vessels that are not prone to extensive motion, digital subtraction angiography (DSA) can be used to remove all background structures from the 2D projection images. TiaVox [37] and Liu et al. [13] successfully reconstruct these sparse structure scenes from only $30$ DSA projections. Liu et al. [13] model blood flow dynamics by decoupling the scene into a static background and dynamic blood vessel foreground. Static and dynamic scene decomposition is one of the approaches to model dynamics with NeRFs, where scenes can be decoupled in a static and dynamic component [26]. However, these DSA scenes do not experience poor vessel visibility, recently highlighted as a significant challenge in sparse-view X-ray angiography reconstruction with NeRFs [7]. Moreover, CA scenes exhibit even more significant dynamics in foreground and background than DSA scenes.
In this work, we decouple the static and dynamic scenes and extend this strategy to address the properties of CA scenes. To the best of our knowledge, NeRF-CA is the first to effectively address the challenges of sparse-views and dynamics of CA data. Inspired by the natural scene decomposition technique D2NeRF [14], we decouple the CA scene into a dynamic coronary artery component and a background component. We uniquely leverage this static and dynamic decomposition by imposing regularization on each learned decoupled component. This regularization is inspired by successful techniques from the natural imaging domain, including structure sparsity [15], scene smoothness [9], and edge occlusion minimization [9]. However, their extension to the CA setting is not straightforward, hence a new methodology is needed. NeRF-CA allows 4D reconstruction from an extremely sparse number of angiogram views without user interaction, which is currently not possible with state-of-the-art methods.